OpenStudy (anonymous):

whats the derivative of ln(3x^2+2)?

OpenStudy (hesan):

1/(3x^2+2) x (6x) =6x/(3x^2+2) This is answer

OpenStudy (campbell_st):

the derivative of a log \[\log(f(x)).... \frac{dy}{dx} =\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

thank you!!!!